Integrated Potato Chips paid a $2.70 per share dividend yesterday. You expect the dividend to grow steadily at a rate of 6% per year.

a.

What is the expected dividend in each of the next 3 years? (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Expected Dividend
  Year 1 $      
  Year 2     
  Year 3     

b.

If the discount rate for the stock is 10%, at what price will the stock sell today? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Current price $

c.

What is the expected stock price 3 years from now? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

  Future price $

d.

If you buy the stock and plan to hold it for 3 years, what payments will you receive? What is the present value of those payments? (Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations. Round your answers to 2 decimal places.)

d.

If you buy the stock and plan to hold it for 3 years, what payments will you receive? What is the present value of those payments? (Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations. Round your answers to 2 decimal places.)

    Year 1     Year 2     Year 3
  Dividend $    $    $   
  Sale of stock         
  Total cash flow $    $    $   
  PV of cash flow $    $    $   
Answer #1

a. Expected dividend payments are:

Dividend in Year 1 = $2.70 * (1 + 6%) = $2.86

Dividend in Year 2 = $2.86 * (1 + 6%) = $3.03

Dividend in Year 3 = $3.03 * (1 + 6%) = $3.22

b. This is an example of constant growth dividend discount model, according to which,

V_{0} = \frac{D_{1}}{r - g}

V0 is the value of share at current year, D1 is the dividend payment in year 1, r is the required rate of return and g is the growth.

Hence,

V0 = 2.86/(10% - 6%) = $71.55

c. Using the same concept as used in part b, expected stock price 3 years from now

V_{3} = \frac{D_{4}}{r - g}

D4 = D3 * (1 + g) = 3.22 * (1 + 6%) = $3.41

V3 = 3.41/(10% - 6%) = $85.22

d.

In this question, we have already ascertained the dividend payments and the future price (at year 3) for this particular share. The only thing here we need to calculate is the PV of all the cashflows.

PV = FV/(1 + r)n (based on TVM concept)

PV (Year 1) = 2.86/(1 + 0.10)1 = $2.60

PV (Year 2) = 3.03/(1 + 0.10)2 = $2.51

PV (Year 3) = 88.43/(1 + 0.10)3 = $66.44

Year 1

Year 2

Year 3

Dividend

$2.86

$3.03

$3.22

Sale of Stock

0

0

$85.22

Total cashflow

$2.86

$3.03

$88.43

PV of cashflow

2.60

2.51

66.44